The area of a rectangle is the product of its length and width, and equals 1062. If the length is 5 more than three times the width, find the width of the rectangles.
Please help me thank you.

Answer by scott8148(6628)

let x=width

"the length is 5 more than three times the width" __ 3x+5

"area of a rectangle is the product of its length and width, and equals 1062"
__ x(3x+5)=1062

3x^2+5x-1062=0 __ factoring __(3x+59)(x-18)=0

3x+59=0 __ x=-59/3 __ negative value not realistic

x-18=0 __ x=18